# STACK: Adversarial Attacks on LLM Safeguard Pipelines

### v1

@@ 1 Introduction
Evaluation: We find black-box STACK bypasses the safeguard models, achieving a 71% ASR on the unambigously harmful query dataset ClearHarm [Hollinsworth et al., 2025]. Transfer STACK achieves a 33% ASR zero-shot, demonstrating attack is possible without any direct interaction with the target pipeline. We conclude with practical recommendations to strengthen defense pipelines.
@@ 2 Related Work
Safeguard Models. Researchers have developed various strategies for enhancing the adversarial robustness of AI systems based on language models. One line of research involves using safeguard models: classifiers that identify harmful content. For example, Wang et al. [2024] developed a transcript classifier approach to prevent models from assisting in bomb-making. Additionally, there has been work to develop more broad safeguard models. Inan et al. [2023] introduced Llama Guard, an LLM-based input-output safeguard designed to classify a range of potential safety risks. Other safeguards include Aegis [Ghosh et al., 2024] and the OpenAI Moderation API [Markov et al., 2022]. In addition to text classifiers, linear activation probes [Alain and Bengio, 2018] have also been applied to classify unsafe model outputs [Ousidhoum et al., 2021], though Bailey et al. [2025] demonstrate that activation monitoring can be circumvented with obfuscated activations. Our work complements existing research by combining safeguard models into a defense-in-depth pipeline and rigorously studying the effectiveness of such systems.
@@
PRP [Mangaokar et al., 2024] attacks models with output classifiers by finding a jailbreak for the classifier and a “propagation prefix” that causes the model to repeat the jailbreak. Concurrent work by Yang et al. [2025a] use a similar idea to PRP to attack vision large language models (VLLMs). Similar to both works, our STACK attack relies on inducing the model to output a jailbreak for the output classifier. However, while PRP could be easily defeated by an input classifier blocking the suspicious query, our attack has been designed to work against models protected with both input and output classifiers via finding universal jailbreaks for each component.
@@ 3 Setting
We conceptualize jailbreaks as being modifications of harmful queries that elicit harmful responses to that query. Formally, we assume there exists a function $H:\Sigma^{*}\times\Sigma^{*}\to[0,1]$ that returns the harmfulness score of a query-response pair $(q,r)$. The attacker can modify their query $q_{0}\mapsto q_{\text{jail}}$ to elicit response $r^{\prime}=p(m)(q_{\text{jail}})$ and then the harmfulness of their response is given by $H(q_{0},r^{\prime})$. That is, the adversary can select the query to generate the response, but the successfulness of their jailbreak attempt will always be evaluated in relation to the original, unobfuscated query $q_{0}$. This criterion ensures that the jailbreak response must be relevant to the attacker’s original query. In practice, we approximate $H$ using an LLM as a judge (Section 3.2), providing it with $q_{0}$ and $r^{\prime}$ in the context window. [^3] We strip output classifier jailbreak attempts from $r^{\prime}$ to avoid accidentally jailbreaking the judge (Appendix E).
@@ 3.1 Motivation
Although our primary focus is misuse, our results also shed light on the efficacy of defense-in-depth pipelines used to guard against other security threats. For example, pipelines may be used to guard against prompt injections in LLM agents [Deng et al., 2025] that would otherwise cause the agent to take actions that violate the developers security policy, e.g., leaking sensitive information [Greshake et al., 2023], or making arbitrary API calls [Pelrine et al., 2024]. Pipelines may also be used to defend LLM overseers that monitor another LLM as part of a control mechanism [Greenblatt et al., 2024].
@@ 3.2 Datasets and Attack Success Rate
StrongREJECT: StrongREJECT provides a general-purpose harmfulness evaluation. It is a dataset of 313 harmful queries across a range of topics including: illegal goods and services; violence; hate, harassment, and discrimination; disinformation and deception; and violence [Souly et al., 2024].
@@
Evaluation: We use the prompted StrongREJECT classifier to compute ASR in both cases. To distinguish between superficially compliant responses, and responses that provide genuinely useful harmful information, only responses considered maximally harmful by the StrongREJECT classifer count as successes for the attacker in our evaluations.
@@ 4.3 Results
We evaluated defense-in-depth pipelines constructed from existing open-weight safeguard models (Figure 3, $x$-axis) against baseline black-box attacks (bar color; Section 4.1) on two datasets (bar shading; Section 3.2). Many of these pipelines perform very poorly against even these simple baseline attacks. For most of these pielines, we concerningly find that they are especially susceptible to attack on the more harmful dataset, ClearHarm (cross-hatch bars).
@@
We find that PAP is the most effective attack, significantly outperforming ReNeLLM and Best-of-N on StrongREJECT. We hypothesise that this is because PAP takes advantage of the dual-use nature of many harmful queries, causing the queries to look benign and the responses to be formatted like ethical responses despite containing harmful information. Indeed, when evaluated on ClearHarm, we find that the attack success rate against few-shot-prompted classifiers “Gwen3” and “Gemma2” plummets to zero.
@@ 6 Limitations and Future Work
Limited classifier efficacy. Although some of the pipelines evaluated significantly improve robustness relative to an undefended model, we find that black-box attacks are effective in most cases even against state-of-the-art open-weight classifiers. This highlights the need for improved classifiers, such as through using larger models, training on more diverse datasets including synthetic data generation, or potentially ensembling multiple different classifiers. We expect our qualitative takeaways to generalize to quantitatively stronger pipelines, however we do suspect the weaknesses of existing classifiers obscures differences in performance between attack methods.
@@ D.2 Confirm hyperparameters
Compute cost. These attacks are possible for a small-scale actor to launch. For the black-box front-to-back attack we spent 507 H100 GPU-days total training all the seeds (costing around \$1000 at current cloud compute prices), and for the white-box transfer attack we spent 15 H100 GPU-days (~\$30) total training all the seeds:
@@ Appendix G Baseline attack prompts
Our implementation of ReNeLLM uses the same rewriting prompts as Ding et al.’s original implementation, except we modified the paraphrase prompt to return one paraphrase (Figure 9) rather than five. We also unintentionally modified the ReNeLLM nesting templates by changing all ellipses from three period characters to a single ellipsis character. Moreover, we unintentionally added an additional backslash before “multicolumn” in the -table-filling nesting template.
@@ Appendix I Black-box Attack Results in Full
To get scores from the pipeline to use for choosing the classifer thresholds (Section 3.3), we first run the input classifier on 4,000 queries from Llama3Jailbreaks and collect the scores. We then generate responses for those queries using the target model Qwen3-14B. We then run the output classifier on the generated responses and collect the scores. We consider all pairs of observed scores $(a,b)$ as possible thresholds for the input and output classifiers respectively if they give an overall Refusal Rate (RR) of less than 15% and are not dominated by another pair in this set. That is, if $\text{RR}(a,b)<\text{RR}(c,d)<=0.15$ and $a=c$ but $b>d$ (i.e. $d$ is more sensitive), then we exclude $(a,b)$ as dominated. We choose thresholds for the input and output classifier from these remaining pairs based on minimizing the absolute difference in individual refusal rates |RR$(f_{q})-\text{RR}(f_{r})\right|| R$.

### v2

@@ 1 Introduction
Evaluation: We find black-box STACK bypasses the safeguard models, achieving a 71% ASR on the unambiguously harmful query dataset ClearHarm [Hollinsworth et al., 2025]. Transfer STACK achieves a 33% ASR zero-shot, demonstrating attack is possible without any direct interaction with the target pipeline. We conclude with practical recommendations to strengthen defense pipelines.
@@ 2 Related Work
Safeguard Models. Researchers have developed various strategies for enhancing the adversarial robustness of AI systems based on language models. One line of research involves using safeguard models: classifiers that identify harmful content. For example, Wang et al. [2024] developed a transcript classifier approach to prevent models from assisting in bomb-making. Additionally, there has been work to develop broader safeguard models. Inan et al. [2023] introduced Llama Guard, an LLM-based input-output safeguard designed to classify a range of potential safety risks. Other safeguards include Aegis [Ghosh et al., 2024] and the OpenAI Moderation API [Markov et al., 2022]. In addition to text classifiers, linear activation probes [Alain and Bengio, 2018] have also been applied to classify unsafe model outputs [Ousidhoum et al., 2021], though Bailey et al. [2025] demonstrate that activation monitoring can be circumvented with obfuscated activations. Our work complements existing research by combining safeguard models into a defense-in-depth pipeline and rigorously studying the effectiveness of such systems.
@@
PRP [Mangaokar et al., 2024] attacks models with output classifiers by finding a jailbreak for the classifier and a “propagation prefix” that causes the model to repeat the jailbreak. Concurrent work by Yang et al. [2025a] uses a similar idea to PRP to attack vision large language models (VLLMs). Similar to both works, our STACK attack relies on inducing the model to output a jailbreak for the output classifier. However, while PRP could be easily defeated by an input classifier blocking the suspicious query, our attack has been designed to work against models protected with both input and output classifiers via finding universal jailbreaks for each component.
@@ 3 Setting
We conceptualize jailbreaks as being modifications of harmful queries that elicit harmful responses to that query. Formally, we assume there exists a function $H:\Sigma^{*}\times\Sigma^{*}\to[0,1]$ that returns the harmfulness score of a query-response pair $(q,r)$. The attacker can modify their query $q_{0}\mapsto q_{\text{jail}}$ to elicit response $r^{\prime}=p(m)(q_{\text{jail}})$ and then the harmfulness of their response is given by $H(q_{0},r^{\prime})$. That is, the adversary can select the query to generate the response, but the success of their jailbreak attempt will always be evaluated in relation to the original, unobfuscated query $q_{0}$. This criterion ensures that the jailbreak response must be relevant to the attacker’s original query. In practice, we approximate $H$ using an LLM as a judge (Section 3.2), providing it with $q_{0}$ and $r^{\prime}$ in the context window. [^3] We strip output classifier jailbreak attempts from $r^{\prime}$ to avoid accidentally jailbreaking the judge (Appendix E).
@@ 3.1 Motivation
Although our primary focus is misuse, our results also shed light on the efficacy of defense-in-depth pipelines used to guard against other security threats. For example, pipelines may be used to guard against prompt injections in LLM agents [Deng et al., 2025] that would otherwise cause the agent to take actions that violate the developer’s security policy, e.g., leaking sensitive information [Greshake et al., 2023], or making arbitrary API calls [Pelrine et al., 2024]. Pipelines may also be used to defend LLM overseers that monitor another LLM as part of a control mechanism [Greenblatt et al., 2024].
@@ 3.2 Datasets and Attack Success Rate
StrongREJECT: StrongREJECT provides a general-purpose harmfulness evaluation. It is a dataset of 313 harmful queries across a range of topics including: illegal goods and services; violence; hate, harassment, and discrimination; and disinformation and deception [Souly et al., 2024].
@@
Evaluation: We use the prompted StrongREJECT classifier to compute ASR in both cases. To distinguish between superficially compliant responses, and responses that provide genuinely useful harmful information, only responses considered maximally harmful by the StrongREJECT classifier count as successes for the attacker in our evaluations.
@@ 4.3 Results
We evaluated defense-in-depth pipelines constructed from existing open-weight safeguard models (Figure 3, $x$-axis) against baseline black-box attacks (bar color; Section 4.1) on two datasets (bar shading; Section 3.2). Many of these pipelines perform very poorly against even these simple baseline attacks. For most of these pipelines, we concerningly find that they are especially susceptible to attack on the more harmful dataset, ClearHarm (cross-hatch bars).
@@
We find that PAP is the most effective attack, significantly outperforming ReNeLLM and Best-of-N on StrongREJECT. We hypothesize that this is because PAP takes advantage of the dual-use nature of many harmful queries, causing the queries to look benign and the responses to be formatted like ethical responses despite containing harmful information. Indeed, when evaluated on ClearHarm, we find that the attack success rate against few-shot-prompted classifiers “Gwen3” and “Gemma2” plummets to zero.
@@ 6 Limitations and Future Work
Limited classifier efficacy. Although some of the pipelines evaluated significantly improve robustness relative to an undefended model, we find that black-box attacks are effective in most cases even against state-of-the-art open-weight classifiers. This highlights the need for improved classifiers, such as through using larger models, training on more diverse datasets including synthetic data generation, or potentially ensembling multiple different classifiers. We expect our qualitative takeaways to generalize to quantitatively stronger pipelines, however we do suspect the weaknesses of existing classifiers obscure differences in performance between attack methods.
@@ D.2 Confirm hyperparameters
Compute cost. These attacks are possible for a small-scale actor to launch. For the black-box front-to-back attack we spent 507 H100 GPU-hours total training all the seeds (costing around \$1000 at current cloud compute prices), and for the white-box transfer attack we spent 15 H100 GPU-hours (~\$30) total training all the seeds:
@@ Appendix G Baseline attack prompts
Our implementation of ReNeLLM uses the same rewriting prompts as Ding et al.’s original implementation, except we modified the paraphrase prompt to return one paraphrase (Figure 9) rather than five. We also unintentionally modified the ReNeLLM templates due to copying them from the paper text rather than the code: for example, this changed all ellipses from three period characters to a single ellipsis character, added an additional backslash before “multicolumn” in the -table-filling nesting template, and removed the braces around the prompt in the code-completion nesting template.
@@ Appendix I Black-box Attack Results in Full
To get scores from the pipeline to use for choosing the classifier thresholds (Section 3.3), we first run the input classifier on 4,000 queries from Llama3Jailbreaks and collect the scores. We then generate responses for those queries using the target model Qwen3-14B. We then run the output classifier on the generated responses and collect the scores. We consider all pairs of observed scores $(a,b)$ as possible thresholds for the input and output classifiers respectively if they give an overall Refusal Rate (RR) of less than 15% and are not dominated by another pair in this set. That is, if $\text{RR}(a,b)<\text{RR}(c,d)<=0.15$ and $a=c$ but $b>d$ (i.e. $d$ is more sensitive), then we exclude $(a,b)$ as dominated. We choose thresholds for the input and output classifier from these remaining pairs based on minimizing the absolute difference in individual refusal rates |RR$(f_{q})-\text{RR}(f_{r})\right|| R$.